# Fast Adaptive Anti-Jamming Channel Access via Deep Q Learning and Coarse-Grained Spectrum Prediction

### v2

@@ 3 Coarse-Grained Spectrum Prediction
Since the legitimate user could not directly observe the jammer’s actions in practical anti-jamming scenarios, inferring the intelligent jammer’s policy using existing opponent modeling approaches becomes challenging. Although the actions of the intelligent jammer are not accessible in practice, the spectrums observed during the current hop could still reflect the behavior of both the legitimate user and the jammers. Therefore, it is possible to employ a convolutional neural network (CNN) model for predicting the spectrum state of the current hop, thereby enhancing the learning capability of the legitimate user’s agent. Since the resolution of the spectrum vector defined in (5) is too fine and thus inefficient for spectrum prediction, we employ the coarse-grained (CG) spectrum with only $M$ samples for each hop.
@@
During the training procedure, the legitimate user collects the spectrum waterfall samples and the ground-truth labels (i.e., the corresponding coarse-grained spectrums) from the real-time interaction between the legitimate user and the wireless adversarial environment. The collected sample-label pair $({\bf{S}}_{n},{\bf{c}}_{n})$ is then stored in the memory $\mathcal{D}_{C}$. However, since the user’s anti-jamming policy and the jammer’s policy are continuously updated through their respective agents during the interactions, the regression model trained on a fixed training dataset may not perform well in a dynamically changing environment. This is because legitimate users with different anti-jamming policies may take entirely different actions when observing the same spectrum waterfall, and the same holds for the intelligent jammer. To address this, we design a dynamically refreshed memory that follows the first-in-first-out (FIFO) principle, storing the latest $|\mathcal{D}_{C}|$ sample-label pairs for training the regression model $F(\cdot;\psi)$.
In each training epoch, a minibatch $\mathcal{B}_{C}$ randomly selected from $\mathcal{D}_{C}$ is fed to the regression model $F(\cdot;\psi)$ for training. The parameters $\psi$ of the regression model could be updated by stochastic gradient descent (SGD), i.e.,
@@
where $\alpha_{C}$ is the learning rate and $L^{C}(\psi)$ is the loss function. To minimize the regression errors, the standard root mean squared error (RMSE) loss is adopted as
@@
denotes the output of the coarse-grained spectrum prediction model, and
@@
denotes the ground-truth label of the sample ${\bf{S}}_{i}$. It is noted that $({\bf{S}}_{i},{\bf{c}}_{i})\in\mathcal{B}_{C}$ is a sample randomly selected from $\mathcal{D}_{C}$, where ${\bf{S}}_{i}$ is the spectrum waterfall composed of several spectrum vectors from previous hops, and the label ${\bf{c}}_{i}$ is the coarse-grained spectrum of the current hop. This enables the network to predict the future coarse-grained spectrum based on the currently observed spectrum waterfall.
@@ 4 Fast Adaptive Anti-jamming Channel Access Approach
As shown in Fig. 5, we construct a Q-function estimation network and a coarse-grained (CG) spectrum prediction network for determining the legitimate transmitter’s channel access action at the beginning of each hop. Specifically, the Q-function estimation network consists of a feature extraction module and an inference module. Meanwhile, the coarse-grained spectrum prediction network is composed of a feature extraction module and a regression forecasting module. It is noted that both the Q-function estimation network and the CG spectrum prediction network take the spectrum waterfall ${\bf{S}}_{n}$ as the input, then output the estimated Q-value and the predicted CG spectrum, respectively. The legitimate user’s anti-jamming action is determined based on the Q-value $Q({\bf{S}}_{n},a_{n}^{u}),\forall a_{n}^{u}\in\mathcal{F}$ and the predicted CG spectrum $\hat{\bf{c}}_{n}$, jointly. The architecture and the training procedure of the proposed fast adaptive anti-jamming channel access approach are discussed in what follows.
@@ 4.1 Q-function Estimation
Since the jammer’s actions are unknown to the legitimate user in practical scenarios, we simplify the proposed anti-jamming MG $\mathcal{G}=\left\{\mathcal{S},\mathcal{A}^{u},\mathcal{A}^{j},\mathcal{P},\mathcal{R}^{u},\mathcal{R}^{j},\gamma\right\}$ to an anti-jamming Markov decision process (MDP), which could be addressed using RL-based methods. The anti-jamming MDP could be formulated by a five-tuple $\left\{\mathcal{S},\mathcal{A}^{u},\mathcal{P},\mathcal{R}^{u},\gamma\right\}$, where $\mathcal{S}$ is the set of environment state, $\mathcal{A}^{u}$ is the legitimate user’s action set, $\mathcal{P}$ denotes the transition function, $\mathcal{R}^{u}$ denotes the legitimate user’s reward function, $\gamma\in(0,1]$ denotes the discount factor.
@@
where $\theta$ denotes the parameters of the Q network and could be optimized by minimizing the loss function $L^{Q}(\theta)$ using gradient descent. The loss function $L^{Q}(\theta)$ is calculated by a random batch of transitions (i.e., $\mathcal{B}_{Q}\subseteq\mathcal{D}_{Q}$) and could be expressed as
@@
is the target Q value of each transition in memory $\mathcal{D}_{Q}$. It is noted that the target network mirrors the Q network, with the exception that its parameters $\theta^{-}$ could be updated by the Q network at a predetermined interval (i.e., $\theta^{-}$ is updated every $N_{u}$ hops).
@@ 4.2 Joint Inference and Decision-Making
To identify the anti-jamming best response policy for the legitimate user, we propose a joint inference and decision-making method aimed at enhancing the state feature representations of the legitimate user’s agent in the proposed anti-jamming scenario. The legitimate user could learn the jammer’s policy features through an auxiliary task (i.e., coarse-grained spectrum prediction) and the jammer’s policy features could serve as a hidden representation for inferring the intelligent jammer’s behavior. Given that the hidden representation encodes the spatial-temporal features of the jammer’s policy $\mu$ and could be learned from the user’s observations over a series of consecutive hops, incorporating the hidden representation into the DQN model could assist in identifying a superior Q-function compared to traditional DRL methods. Similarly, the hidden features extracted by the DQN model could also enhance the learning capability of the coarse-grained spectrum prediction model.
The network architecture of the joint inference and decision-making model is illustrated in Fig. 6. The model is composed of a feature extraction module, a coarse-grained spectrum prediction module, a Q-function estimation module, and a joint decision function. The feature extraction module employs a CNN with two Conv layers to extract features from the currently observed spectrum waterfall. The feature extracted from the spectrum waterfall ${\bf{S}}_{n}$ is denoted as $h_{n}^{e}$.
@@
It is important to note that the feature extraction module is shared between the coarse-grained spectrum prediction module and the Q-function estimation module. Subsequently, these two modules take $h_{n}^{e}$ as their common input, respectively. Both the coarse-grained spectrum prediction module and the Q-function estimation module consist of several FC layers. The coarse-grained spectrum prediction module is trained to infer the CG spectrum during the current hop, while the Q-function estimation module is trained to approximate the optimal Q-function $Q^{*}({\bf{S}},a)$. The two modules operate independently to extract two distinct types of features (i.e., $h_{n}^{C}$ and $h_{n}^{Q}$) from $h_{n}^{e}$ using an FC layer, and these features are concatenated to construct the hidden representation $h_{n}=\left(h_{n}^{Q};h_{n}^{C}\right)$. After that, $h_{n}$ is processed by two FC layers to generate the predicted coarse-grained spectrum during the current hop (i.e., $\hat{\bf{c}}_{n}$) within the coarse-grained spectrum prediction module and the estimated Q-values (i.e., $Q({\bf{S}}_{n},a^{u}),\ \forall a^{u}\in\mathcal{A}^{u}$) within the Q-function estimation module, respectively. Ultimately, the legitimate user’s anti-jamming channel access action is determined by
@@
where $\hat{\bf{c}}_{n}[a^{u}]$ denotes the $a^{u}$-th element of the predicted coarse-grained spectrum $\hat{\bf{c}}_{n}$. It is noted that this decision-making function is merely one of the viable functions, which could be further refined to suit the specific anti-jamming scenario.
@@
The training procedure of the proposed joint inference and decision-making model is detailed in Algorithm 1. We incorporate two distinct loss terms (i.e., $L^{C}(\psi)$ defined in (22) and $L^{Q}(\theta)$ defined in (29)) to construct an aggregated loss function $L(\theta,\psi)$ for training the proposed anti-jamming model, jointly. The aggregated loss $L(\theta,\psi)$ is defined as
$L(\theta,\psi)=\lambda L^{Q}(\theta)+L^{C}(\psi),$ (33)
where $L^{C}(\psi)$ is the regression forecasting loss for coarse-grained spectrum prediction, $L^{Q}(\theta)$ is the standard DQN loss function, $\lambda=\frac{1}{\sqrt{L^{C}(\psi)}}$ serves as the adaptive scale factor of $L^{Q}(\theta)$, which could adaptively adjust the scaling of $L^{Q}(\theta)$ during different phases of the training process.
@@ 5.1 Simulation Setup
Simulations are presented to validate the superiority of the proposed anti-jamming approach, with parameter settings referenced from Wang et al. (2020); Liu et al. (2018). In experiments, the frequency band from 0MHz to 20MHz (i.e., the total bandwidth $B=20$MHz) is divided into $M=10$ non-overlapping channels, with the bandwidth of each channel being $b=2$MHz. The legitimate receiver and the intelligent jammer could perform spectrum sensing actions every $\Delta t=1$ms with $\Delta f=100$kHz, retaining the spectrum data for $T=200$ms. The number of samples in each spectrum vector is $N_{F}=B/\Delta f=200$ and the length of historical data is $N_{T}=T/\Delta t=200$. Therefore, the size of the spectrum waterfall ${\bf{S}}_{n}$ is $N_{F}\times N_{T}=200\times 200$. The fixed-mode jammer could launch traditional jamming attacks with a fixed jamming mode, while the intelligent jammer could attack $N_{I}=3$ consecutive channels in each time slot, with an update step 10 times that of the user. Additionally, the hyper-parameters of the intelligent jammer’s agent are set according to Li et al. (2023a). The bandwidth of each jamming tone is 2MHz, and its power is set to 50dBm.
The legitimate signal has a bandwidth of 2MHz, and its working frequency is allowed to hop every $T_{h}=10$ms. The transmission power of the legitimate transmitter is 30dBm, and the legitimate receiver’s demodulation threshold is set to $\beta_{\text{th}}=0$dB. The legitimate signal is shaped with a root-raised cosine pulse shaping filter with a roll-off factor of 0.5. The hyper-parameters in Algorithm 1 are set as follows: learning rate is $\alpha_{Q}=\alpha_{C}=1\times 10^{-4}$, minibatch size is $|\mathcal{B}_{Q}|=|\mathcal{B}_{C}|=64$, size of memory $\mathcal{D}_{Q}$ is $|\mathcal{D}_{Q}|=1000$, size of memory $\mathcal{D}_{C}$ is $|\mathcal{D}_{C}|=256$, update frequency of target Q network is $N_{u}=1000$, discount factor is $\gamma=0.1$. The Q-function estimation network and the coarse-grained spectrum prediction network share the same CNN architecture, which is detailed in Table 1.
@@
FC1 $32\times 50\times 50$ $512$ ReLU
FC2 $512\times 2$ $256$ ReLU
FC3 $256$ $10$

### v3

@@ 3 Coarse-Grained Spectrum Prediction
Since the legitimate user could not directly observe the jammer’s actions in practical anti-jamming scenarios, inferring the intelligent jammer’s policy using existing opponent modeling approaches becomes challenging. Although the actions of the intelligent jammer are not accessible in practice, the spectra observed during the current hop could still reflect the behavior of both the legitimate user and the jammers. Therefore, it is possible to employ a convolutional neural network (CNN) model for predicting the spectrum state of the current hop, thereby enhancing the learning capability of the legitimate user’s agent. Since the resolution of the spectrum vector defined in eq. (5) is too fine and thus inefficient for spectrum prediction, we employ the coarse-grained (CG) spectrum with only $M$ samples for each hop.
@@
During the training procedure, the legitimate user collects the spectrum waterfall samples and the ground-truth labels (i.e., the corresponding coarse-grained spectra) from the real-time interaction between the legitimate user and the wireless adversarial environment. The collected sample-label pair $({\bf{S}}_{n},{\bf{c}}_{n})$ is then stored in the memory $\mathcal{D}_{C}$. However, since the user’s anti-jamming policy and the jammer’s policy are continuously updated through their respective agents during the interactions, the regression model trained on a fixed training dataset may not perform well in a dynamically changing environment. This is because legitimate users with varying anti-jamming policies may take entirely different actions when observing the same spectrum waterfall, and the same holds for an intelligent jammer. To address the non-stationarity characteristic of the environment and enable the model to adapt to recent changes in the behavior of both the legitimate user and the jammers, the spectrum prediction model is not trained on the entire accumulated dataset. Instead, a dynamically refreshed memory is employed, which follows the first-in-first-out (FIFO) principle and stores the most recent $|\mathcal{D}_{C}|$ sample-label pairs for training the regression model $F(\cdot;\psi)$.
In each training epoch, the memory $\mathcal{D}_{C}$ is fed to the regression model $F(\cdot;\psi)$ for training, where one epoch corresponds to a complete pass through the entire training dataset. The parameters $\psi$ of the regression model could be updated by stochastic gradient descent (SGD), i.e.,
@@
where $\alpha_{C}$ is the learning rate and $L^{C}(\psi)$ is the loss function. To minimize the expected regression error, the standard root mean squared error (RMSE) loss is adopted as
@@
denotes the output of the CG spectrum prediction model, and
@@
denotes the ground-truth label of the sample ${\bf{S}}_{i}$. It is noted that $({\bf{S}}_{i},{\bf{c}}_{i})\in\mathcal{D}_{C}$ is a sample in $\mathcal{D}_{C}$, where ${\bf{S}}_{i}$ is the spectrum waterfall composed of several spectrum vectors from previous hops, and ${\bf{c}}_{i}$ is the coarse-grained spectrum of the current hop. This enables the network to predict the future coarse-grained spectrum based on the currently observed spectrum waterfall. In order to enable the proposed coarse-grained spectrum prediction model to adapt to the non-stationary nature of the environment, we do not use all the accumulated data, as is typically done in conventional supervised learning, for model training. Instead, at each training iteration, only the most recent $|\mathcal{D}_{C}|$ samples are employed to update the parameters of the coarse-grained spectrum prediction model. This strategy ensures that the prediction model remains sensitive to the latest changes in the environment, including the evolving transmission behavior of legitimate users and adversarial jamming strategies. By focusing on recent samples, the model could better track the dynamic spectrum patterns and maintain high prediction accuracy in non-stationary scenarios.
@@ 4 Fast Adaptive Anti-jamming Channel Access Approach
As shown in Fig. 6, we construct a Q-function estimation network and a CG spectrum prediction network for determining the legitimate transmitter’s channel access action at the beginning of each hop. Specifically, the Q-function estimation network consists of a feature extraction module and an inference module. Meanwhile, the coarse-grained spectrum prediction network is composed of a feature extraction module and a regression forecasting module. It is noted that both the Q-function estimation network and the CG spectrum prediction network take the spectrum waterfall ${\bf{S}}_{n}$ as the input, then output the estimated Q-value and the predicted CG spectrum, respectively. The legitimate user’s anti-jamming action is determined based on the Q-value $Q({\bf{S}}_{n},a_{n}^{u}),\forall a_{n}^{u}\in\mathcal{F}$ and the predicted CG spectrum $\hat{\bf{c}}_{n}$, jointly. The architecture and the training procedure of the proposed fast adaptive anti-jamming channel access approach are discussed in what follows.
@@ 4.1 Q-function Estimation
Since the jammer’s actions are unknown to the legitimate user in practical scenarios, we simplify the proposed anti-jamming MG $\mathcal{G}=\left\{\mathcal{S},\mathcal{A}^{u},\mathcal{A}^{j},\mathcal{P},\mathcal{R}^{u},\mathcal{R}^{j},\gamma\right\}$ to an anti-jamming MDP, which could be addressed using RL-based methods. The anti-jamming MDP could be formulated by a five-tuple $\left\{\mathcal{S},\mathcal{A}^{u},\mathcal{P},\mathcal{R}^{u},\gamma\right\}$, where $\mathcal{S}$ is the set of environment state, $\mathcal{A}^{u}$ is the legitimate user’s action set, $\mathcal{P}$ denotes the transition function, $\mathcal{R}^{u}$ denotes the legitimate user’s reward function, $\gamma\in(0,1]$ denotes the discount factor.
@@
where $\theta$ denotes the parameters of the Q network and could be optimized by minimizing the loss function $L^{Q}(\theta)$ using gradient descent. $L^{Q}(\theta)$ is calculated by a random batch of transitions (i.e., $\mathcal{B}_{Q}\subseteq\mathcal{D}_{Q}$) and could be expressed as
@@
is the target Q value of each transition in $\mathcal{D}_{Q}$. It is noted that the target network mirrors the Q network, with the exception that its parameters $\theta^{-}$ could be updated by the Q network at a predetermined interval (i.e., $\theta^{-}$ is updated every $N_{u}$ hops).
@@ 4.2 Joint Inference and Decision-Making
To identify the anti-jamming BR policy for the legitimate user, we propose a joint inference and decision-making method aimed at enhancing the state feature representations of the legitimate user’s agent in the proposed anti-jamming scenario. The legitimate user could learn the jammer’s policy features through an auxiliary task (i.e., CG spectrum prediction) and the jammer’s policy features could serve as a hidden representation for inferring the intelligent jammer’s behavior. Given that the hidden representation encodes the spatial-temporal features of the jammer’s policy $\mu$ and could be learned from the user’s observations over a series of consecutive hops, incorporating the hidden representation into the DQN model could assist in identifying a superior Q-function compared to traditional DRL methods. Similarly, the hidden features extracted by the DQN model could also enhance the learning capability of the coarse-grained spectrum prediction model.
We design a dual-branch CNN architecture to jointly perform coarse-grained spectrum prediction and Q-value estimation, as illustrated in Fig. 7. The input spectrum waterfall ${\bf{S}}_{n}$ is initially processed by two shared convolutional layers to extract time-frequency features, resulting in a latent representation, denoted as $h_{n}^{e}$. This representation is then fed into two parallel branches: one for spectrum prediction and the other for decision evaluation. In the prediction branch, $h_{n}^{e}$ is processed by an FC layer to produce $h_{n}^{C}$. In the evaluation branch, $h_{n}^{e}$ is similarly transformed into $h_{n}^{Q}$ via an FC layer. Then the two representations $h_{n}^{Q}$ and $h_{n}^{C}$ are concatenated to form a joint feature representation $h_{n}$, which captures both prediction and evaluation cues. This shared representation is then separately fed into task-specific FC layers in the two branches to generate the coarse-grained spectrum prediction and estimate the Q-value, respectively. Such a design facilitates effective feature sharing while maintaining task-specific outputs.
@@
It is important to note that the feature extraction module is shared between the coarse-grained spectrum prediction module and the Q-function estimation module. Subsequently, these two modules take $h_{n}^{e}$ as their common input, respectively. Both the coarse-grained spectrum prediction module and the Q-function estimation module consist of several FC layers. The coarse-grained spectrum prediction module is trained to infer the CG spectrum during the current hop, while the Q-function estimation module is trained to approximate the optimal Q-function. The two modules operate independently to extract two distinct types of features (i.e., $h_{n}^{C}$ and $h_{n}^{Q}$) from $h_{n}^{e}$ using an FC layer, and these features are concatenated to construct the hidden representation $h_{n}=\left(h_{n}^{Q};h_{n}^{C}\right)$. After that, $h_{n}$ is processed by two FC layers to generate the predicted coarse-grained spectrum during the current hop (i.e., $\hat{\bf{c}}_{n}$) within the coarse-grained spectrum prediction module and the estimated Q-values (i.e., $Q({\bf{S}}_{n},a^{u}),\ \forall a^{u}\in\mathcal{A}^{u}$) within the Q-function estimation module, respectively. Ultimately, the legitimate user’s anti-jamming channel access action is determined by
@@
where $Q({\bf{S}}_{n},a^{u})$ denotes the predicted Q-value of channel $a^{u}$ under the current state ${\bf{S}}_{n}$, and $\hat{\bf{c}}_{n}[a^{u}]$ represents the predicted coarse-grained spectrum sample (in dBm) for channel $a^{u}$. This function is intended to prioritize channels that simultaneously offer high expected long-term reward and low predicted spectrum sample value. The denominator converts the spectrum sample value from dBm to mW, thereby aligning it with the magnitude of the Q-value in the numerator and ensuring that both components contribute comparably to the decision-making process. It is noted that this decision-making function is merely one of the viable functions, which could be further refined to suit the specific anti-jamming scenario.
It is noted that, in the proposed joint decision mechanism, the coarse-grained spectrum prediction results are not used as a standalone decision but instead serve as an input to the joint decision module. During the early training stages, these predictions help improve the quality of action selection over random exploration. As a result, the agent is more likely to obtain positive reward samples early on, which in turn accelerates the rate of convergence of the Q-network and enhances the anti-jamming performance. This design improves the exploration efficiency of the DRL framework without compromising its learning capability.
The training procedure for the proposed joint inference and decision-making model is detailed in Algorithm 1, where $L^{C}_{\text{th}}$ serves as the threshold to balance exploration and exploitation for the anti-jamming agent. We incorporate two distinct loss terms (i.e., $L^{C}(\psi)$ defined in eq. (22) and $L^{Q}(\theta)$ defined in eq. (29)) to construct an aggregated loss function $L(\theta,\psi)$ for training the proposed anti-jamming model, jointly. The aggregated loss $L(\theta,\psi)$ is defined as
$L(\theta,\psi)=\lambda L^{Q}(\theta)+L^{C}(\psi),$ (33)
where $L^{C}(\psi)$ is the regression forecasting loss for coarse-grained spectrum prediction, $L^{Q}(\theta)$ is the standard DQN loss function, $\lambda=\frac{1}{\sqrt{L^{C}(\psi)}}$ serves as the adaptive scale factor of $L^{Q}(\theta)$, which could adaptively adjust the scaling of $L^{Q}(\theta)$ during different phases of the training process. Unlike conventional fixed-weight strategies, our framework adaptively adjusts $\lambda$ during the training process. This allows the model to prioritize prediction accuracy during early training and gradually emphasize value estimation as learning progresses. Such dynamic re-weighting enhances training stability and fosters better synergy between the two learning objectives.
@@ 5.1 Simulation Setup
Simulations are presented to validate the superiority of the proposed anti-jamming approach, with parameter settings referenced from Wang et al. (2020); Liu et al. (2018). In experiments, the frequency band from 0MHz to 20MHz (i.e., the total bandwidth B=20B=20MHz) is divided into $M=10$ non-overlapping channels, with the bandwidth of each channel being b=2b=2MHz. The path loss exponent is set to $\alpha_{d}=1$, and the instantaneous fading coefficient $\xi_{k}$ follows an exponential distribution with a mean of 1 Wu et al. (2013). The legitimate receiver and the intelligent jammer could perform spectrum sensing actions every $\Delta t=1$ms with $\Delta f=100$kHz, retaining the spectrum data for T=200T=200ms. The number of samples in each spectrum vector is $N_{F}=B/\Delta f=200$ and the length of historical data is $N_{T}=T/\Delta t=200$. Therefore, the size of the spectrum waterfall ${\bf{S}}_{n}$ is $N_{F}\times N_{T}=200\times 200$. The fixed-mode jammer could launch traditional jamming attacks with a predetermined jamming mode, while the intelligent jammer could attack $N_{I}=3$ consecutive channels in each time slot, with an update step 10 times that of the legitimate user Li et al. (2021, 2023c). Additionally, the hyper-parameters of the intelligent jammer’s agent are set according to Li et al. (2023c). The bandwidth of each jamming tone is 2MHz, and its power is set to 50dBm.
The legitimate signal has a bandwidth of 2MHz, and its working frequency is allowed to hop every $T_{h}=10$ms. The transmission power of the legitimate transmitter is 30dBm, and the legitimate receiver’s demodulation threshold is set to $\beta_{\text{th}}=0$dB. The legitimate signal is shaped with a root-raised cosine pulse shaping filter with a roll-off factor of 0.5. The hyper-parameters in Algorithm 1 are set as follows: learning rate is $\alpha_{Q}=\alpha_{C}=1\times 10^{-4}$, minibatch size is $|\mathcal{B}_{Q}|=64$, size of memory $\mathcal{D}_{Q}$ is $|\mathcal{D}_{Q}|=1000$, size of memory $\mathcal{D}_{C}$ is $|\mathcal{D}_{C}|=64$, update frequency of target Q network is $N_{u}=1000$, discount factor is $\gamma=0.1$, loss threshold is $L^{C}_{\text{th}}=10$. The Q-function estimation network and the coarse-grained spectrum prediction network share the same CNN architecture, which is detailed in Table 1. During the training procedure, one episode consists of 100 hops, with the parameters of both the coarse-grained spectrum prediction network and the Q-function estimation network updated at the end of each hop.
@@
FC1 $32\times 50\times 50$ $512$ - ReLU
FC2 $512\times 2$ $256$ - ReLU
FC3 $256$ $10$ - -